OpenStudy (anonymous):

please show steps: Solve: log base 5 3-1/2 log base 5 x=log base 5 1/2

OpenStudy (dumbcow):

use log identities: http://en.wikipedia.org/wiki/Logarithm#Product.2C_quotient.2C_power_and_root \[\log_{5}3 - \frac{1}{2}\log_{5} x = \log_{5} \frac{1}{2} \] \[\log_{5}3 - \log_{5} \sqrt{x} = \log_{5} \frac{1}{2}\] \[\log_{5} \frac{3}{\sqrt{x}} = \log_{5} \frac{1}{2}\] \[\frac{3}{\sqrt{x}} = \frac{1}{2}\] \[\sqrt{x} = 6\] \[x = 36\]